(– 43) × (– 99) + 43 is equal to
(a) 4300 (b) – 4300 (c) 4257 (d) – 4214

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

Option (a) is correct here
By BODMAS rule,
(–43) × (–99) + 43 = [(–43) × (–99)] + 43
= 4257 + 43
= 4300

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
Applying BODMAS,
(-43)×(-99) + 43 
= [(-43)×(-99)] + 43
= 4357 + 43
= 4300
 
Hope it helps!
Thank you